SourMunchkin7806:

Find the inverse of the function. f(x) = 4x - 6 Not a one-to-one function f-1(x) = quantity x plus six divided by four f-1(x) = x divided by four plus six f-1(x) = quantity x minus six divided by four

SourMunchkin7806:

@Hero

SourMunchkin7806:

@Vocaloid

SourMunchkin7806:

@Isry

SourMunchkin7806:

@Hero

Hero:

First change f(x) to y, then swap the x and y. Solve for y. Finally change to f^(-1)(x)

SourMunchkin7806:

So would this be D?

Hero:

Have you solved it yet or are you just guessing?

SourMunchkin7806:

well im doing this in my head so mainly guess...but i also suck terribly at this so i dont really know how to solve for y here

Hero:

If you change f(x) to y what will you have afterwards? Can you type it here please?

SourMunchkin7806:

Yea then we would have y = 4x - 6 right?

Hero:

Great, now swap the positions of x and y. Post the result here.

SourMunchkin7806:

ok so now its x = 4y - 6

Hero:

Excellent. Now add 6 to both sides. Post the result here.

SourMunchkin7806:

x = 4y?

Hero:

Did you add 6 to BOTH sides?

SourMunchkin7806:

yes according to what i was taught...im guessing what i was taught was wrong

Hero:

Okay, if you start with x = 4y - 6 then add 6 to both sides, what would that look like? Can you show me your process?

SourMunchkin7806:

so 6 + x = 4y - 6 + 6

Hero:

Great. Now simplify that. Post it here.

SourMunchkin7806:

x6 = 4y?

SourMunchkin7806:

sorry man math is not my best subject

Hero:

Explain why you remove the PLUS on the left side.

SourMunchkin7806:

im not sure thats just how all the examples in my lessons show it so i thought it was right...would it be 6 + x = 4y then?

Hero:

Yes, correct. It is proper to write the variable first. In that case, you would have x + 6 = 4y. Now the last step is to do what? Any idea what the last step might be?

SourMunchkin7806:

no...

SourMunchkin7806:

wait do we subtract x from each side?

Hero:

Divide both sides by 4. Remember, the goal was to isolate y because once you did that you would have found the inverse of that. Go ahead and divide both sides by 4. Let me know what you get.

SourMunchkin7806:

i dont know...but i think i know the answer to the question...as i think i just have to find the equation

SourMunchkin7806:

can you help me with one more question maybe...

Hero:

Are you telling me that with one step remaining, you are unable to figure out how to express the correct answer? Can you at least make the attempt please. I'd prefer if you used the drawing tool for this. Thanks.

SourMunchkin7806:

well the end i came up with is 

SourMunchkin7806:

sorry on a laptop

SourMunchkin7806:

so the right side i got y as the 4 kinda cancels out and the left side i cant solve

Hero:

Yes, that is correct. So using \(\LaTeX\) to express the correct result we get \(y = \dfrac{x + 6}{4}\) but remember we replace \(y\) with \(f^{-1}(x)\) to get \(f^{-1}(x) = \dfrac{x + 6}{4}\)

SourMunchkin7806:

awesome thank you so much

SourMunchkin7806:

as i said i kinda have on more if you can help me with that one

SourMunchkin7806:

one*

Hero:

Are you sure you're not able to finish it up on your own?. Go ahead, post it, then post all your steps. Make sure to tag me AFTER you have finished with your attempt. Deal?

SourMunchkin7806:

no like i dont even know where to start im supposed to explain how two things are inverses f(x)= 4 over x and g(x) = 4 over x and i dont know where i am supposed to start with that like at all

Hero:

Post it as a separate post my friend.